OpenStudy (anonymous):

how do I simplify square root of 540?

OpenStudy (anonymous):

\[\sqrt{540}\]

OpenStudy (jdoe0001):

 \(\bf \sqrt{540}\\ \quad \\ \quad \\ \color{blue}{540 = 2 \cdot 2\cdot 3\cdot 3\cdot 3\cdot 3\cdot 5\implies 2^2 \cdot 3^2 \cdot 3 \cdot 5}\\ \quad \\ \sqrt{540}\implies \sqrt{2^2 \cdot3^2 \cdot 3 \cdot 5}\)